OpenStudy (anonymous):

Tell whether the equation -4x + 2y = -2 represents a direct variation. If so, identify the constant of variation. Direct variation; k = ½ Direct variation; k = 2 Direct variation; k = -2 Not a direct variation

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Ddos_Dragon

Nnesha (nnesha):

what's the general equation for direct variation ?

OpenStudy (anonymous):

Hmmm... I dunno.

OpenStudy (anonymous):

\[\frac{ y }{ x } = k\] ?

Nnesha (nnesha):

i helped you yesterday i asked same question.. :P

Nnesha (nnesha):

that's correct but you should write it as `y=kx`

Nnesha (nnesha):

where k is the constant of variation now solve the given equation for y

Nnesha (nnesha):

the direct variation equation is in y= form so you should solve the original equation for

Nnesha (nnesha):

for y*

OpenStudy (anonymous):

-4x + 2y = -2 -4 / -4x + 2y = -2 / -4 x + 2y = 0.5 x + 2 / 2y = 0.5 /2 x + y = 0.25 ?

OpenStudy (anonymous):

I'm unsure if I did what you asked for.

Nnesha (nnesha):

hmm good job for trying :=)) but hmm sorry that's wrong you \[\huge\rm -4x+2y=-2\] in order to solve for y first you should move the x term to the right side (not just the coefficient *x term* whihc is -4x)

OpenStudy (anonymous):

So.. to do that we'd need to add a -4 and a x to cancel them out on the left side? 2y = -2 + -4/x ?

Nnesha (nnesha):

add -4x both side remember this is one term `-4x`

Nnesha (nnesha):

 you can't add -2 with 4x they are not like terms so just leave 4x at right side